If 36.5 mol of an ideal gas occupies 66.5 L at 473 K, what is the pressure of the gas? atm pressure: TOOLS x10
Answer #1

Given:
V = 66.5 L
n = 36.5 mol
T = 473.0 K

use:
P * V = n*R*T
P * 66.5 L = 36.5 mol* 0.08206 atm.L/mol.K * 473 K
P = 21.3 atm
Answer: 21.3 atm